# AWARE: Audio Watermarking via Adversarial Resistance to Edits

### v1

@@ Abstract
Prevailing practice in learning-based audio watermarking is to pursue robustness by expanding the set of simulated distortions during training. However, such surrogates are narrow and prone to overfitting. This paper presents AWARE (Audio Watermarking with Adversarial Resistance to Edits), an alternative approach that avoids reliance on attack-simulation stacks and handcrafted differentiable distortions. Embedding is obtained via adversarial optimization in the time–frequency domain under a level-proportional perceptual budget. Detection employs a time–order–agnostic detector with a Bitwise Readout Head (BRH) that aggregates temporal evidence into one score per watermark bit, enabling reliable watermark decoding even under desynchronization and temporal cuts. Empirically, AWARE attains high audio quality and speech intelligibility (PESQ/STOI) and consistently low BER across various audio edits, often surpassing representative state-of-the-art learning-based audio watermarking systems.
@@ 1 Introduction
Digital watermarking experienced its first major wave of research activity in the 1990s alongside the rapid proliferation of the Internet. Early systems were primarily designed for copyright protection and digital rights management (DRM), with the seminal work of Cox et al. introducing spread-spectrum principles to watermarking and setting the agenda for robustness-focused design [1]. Subsequent developments broadened the methodological toolbox with techniques such as quantization index modulation (QIM) [2] and patchwork-style techniques [3]. While initial approaches were conceived as modality-agnostic and applicable across multimedia, the field soon bifurcated into image- and audio-specific lines of work [4, 5], each exploiting modality characteristics to improve embedding efficiency and detection reliability. The dominant use case remained copyright protection, driven by the rise of large-scale online content distribution and associated piracy.
@@
Despite their impact, traditional watermarking techniques (particularly in audio) face persistent limitations. Beyond limited robustness to classical signal processing, two challenges loom large: desynchronization and waveform cuts. Systems often include dedicated synchronization codes to address time-scale modifications, resampling drift, cropping, and jitter. However, reliably detecting synchronization markers is nearly as difficult as extracting the watermark itself, and thus inherits similar failure modes under distortion. Moreover, many legacy designs embed watermark bits within a single frame or a narrow group of frames, yielding limited temporal redundancy and weak fragment-level detectability. Missing or re-ordered frames, or partial content removal, can therefore break the decoding process and make the watermark hard or impossible to reconstruct.
@@
The advent of modern generative AI has precipitated a renaissance in digital audio watermarking. High-fidelity synthesis models such as: GANs [6] and diffusion models [7] enable convincing audio and audiovisual “deepfakes” at scale. The risks span reputational harm, fraud, misinformation, and weakened evidence reliability. In response, watermarking has re-emerged as a practical mechanism to label both synthetic and authentic content to support provenance, traceability, and downstream moderation. Accordingly, policy frameworks increasingly cite watermarking among key techniques for AI transparency and content provenance [8].
Concurrently, the community has begun to “fight fire with fire,” developing end-to-end deep learning (DL) audio watermarking systems. Leading this line of research, RobustDNN [9] defined the basic blueprint, after which WavMark [10] and AudioSeal [11] introduced meaningful improvements. Nevertheless, contemporary benchmark studies still indicate unresolved limitations, with a clear room for progress [12, 13].
Critically, watermark decoding for audio under temporal cuts remains underexplored, aside from zero-bit approaches [14] that have limited practical scope. Unlike images, where spatial cropping typically retains substantial context, audio pipelines frequently yield spliced content: selective cuts, concatenation of short segments from different sources, or montage-like edits. Such edits can sound natural to human listeners, yet disrupt global synchronization and erase large portions of the embedded message. Practical deployments therefore require watermarking that survives cuts and splicing. Meanwhile, high-fidelity voice cloning introduces a new problem that challenges established provenance mechanisms.
A general trend in deep learning has been to port high-performing architectures from other domains, most notably computer vision, by “stacking layers” from CNN/Transformer backbones. This pattern has also influenced DL-based audio watermarking. Classic signal-processing–driven designs, by contrast, were intentionally constrained: each design choice (domain selection, embedding operator, synchronization mechanism, detection statistic) was motivated by the signal model and anticipated threat model.
While modern DL-based approaches have accelerated progress, they often overlook the unique nature of watermark detection in audio. Watermark detection differs fundamentally from object/keyword detection or semantic classification. The target is not a semantic entity localized in space or time. It is a weak, distributed pattern, encoding bits that must be sequence-consistent under time-warping and cutting. The decoder must aggregate evidence over time, maintain or recover alignment, and ultimately produce a bitstream (with reliability scores).
To address these gaps, this paper introduces AWARE (Audio Watermarking with Adversarial Resistance to Edits): an adversarial watermarking procedure "aware" of auditory perception and audio signal structure. It employs an adversarial embedding procedure under a level-proportional perceptual budget and a time–order–agnostic detector with a Bitwise Readout Head that aggregates temporal evidence into per-bit scores. Comprehensive experiments show high perceptual quality and intelligibility with consistently low BER under diverse edits, often surpassing state-of-the-art baselines and yielding a more stable BER profile across conditions, including desynchronization and temporal cuts.
@@ 2 Background
In the standard adversarial setting, we start with a clean input $x$, its label $y$, a model $f$, and a task-specific loss function $\mathcal{L}$. The objective is to find a small, norm-bounded perturbation $\Delta$ that maximizes the loss $\mathcal{L}$:
$\max_{\lVert\Delta\rVert_{p}\leq\epsilon}\ \mathcal{L}\big(f(x+\Delta),\,y\big).$ (1)
Intuitively, we seek the smallest change (constrained by $\|\Delta\|_{p}\leq\varepsilon$) that maximizes the degradation of the model’s performance on $x$. The choice of $p$ (e.g., $p=\infty$ or $2$) and radius $\varepsilon$ encodes the allowable perturbation “budget”.
Adversarial perturbations have proved effective across diverse tasks within the audio modality [15, 16, 17], predominantly as tools to break models rather than to enforce desired behaviors. Li et al. [18] argue that watermarking is essentially a signal perturbation optimized to steer the detector’s outputs. Hence, they introduce adversarial shallow watermarking for images, pairing a frozen detector with an embedding procedure that adjusts the carrier until a randomly initialized detector yields the desired watermark bits. Two insights are particularly relevant to our setting: (i) detector architectures should remain shallow to improve out-of-distribution generalization and resilience to unseen attacks, and (ii) distortion-simulation stacks can be simplified or eliminated, avoiding reliance on large suites of handcrafted differentiable attacks.
Learning-based watermarking often gains robustness by inserting differentiable “attack layers” into training. However, this strategy faces practical limits: the space of plausible attacks is vast (combinatorial in type and composition), so exhaustive coverage during training is infeasible. Moreover, Li et al. observe that DL-based systems tend to overfit to the attack set seen during training and generalize poorly when novel perturbations are introduced.
This adversarial approach has another advantage that it enables rapid evaluation of different architectural choices under a broad set of perturbations by avoiding lengthy supervised training. The resulting insights can later inform the design of learning-based watermarking systems (with or without attack layers).
Naively adapting the Li et al. approach to audio waveforms/spectrograms inherits the same issues seen in many DL-based audio watermarking systems under cuts and desynchronization, as the method was tailored to images. Audio watermark decoders require architectural mechanisms that are intrinsically robust to common edits and temporal misalignments. If the architecture is not inherently robust, no amount of augmentation or attack-layer engineering will make training reliably effective.
@@ 3 AWARE: Method
Watermark Embedding. Embedding is carried out in the time-frequency (TF) domain, a standard and effective setting for audio watermarking. The complete procedure is outlined in Algorithm 1.
@@
Let $x\in\mathbb{R}^{T}$ be a waveform and let $\lvert\mathrm{STFT}(x)\rvert\in\mathbb{R}_{\geq 0}^{F\times U}$ denote its short-time Fourier transform (STFT) magnitude, with frequency bins $f\in[0,\dots,F-1]$ and time frames $u\in[0,\dots,U-1]$. We restrict perturbations to an audible midband $\mathcal{F}=\{f:f_{\ell}\leq f\leq f_{h}\}$ with $f_{\ell}=500~\text{Hz}$ and $f_{h}=4000~\text{Hz}$ to avoid removal by low/high-pass filters.
Phase-domain watermark embedding is avoided due to human hearing being largely insensitive to changes in phase, which allows for low-audibility phase manipulations that effectively erase the mark. Accordingly, we modify magnitudes under perceptual constraints and preserve the original phase for reconstruction (iSTFT).
We denote $w$ as a watermark of $N$ bits, and $D$ represents a randomly initialized detector with frozen weights that maps a TF magnitude representation of $x$ to $\left(-1,+1\right)^{N}$. Watermark bits are encoded antipodally as $\tilde{w}\in\{-1,+1\}^{N}$. This centers targets at zero, yields balanced gradients, and pairs naturally with margin-based objectives and sign decoding, unlike $\{0,1\}$ coding which pushes scores towards probability bounds and can hinder optimization.
@@
Rather than imposing a single global norm budget on $\Delta$, we use a per-bin, level-proportional budget that allows larger changes where the signal is louder and smaller changes in quiet regions, consistent with basic psychoacoustics. Let $\tau_{\text{dB}}>0$ be a tolerance parameter (in dB). Its linear amplitude factor is $\eta\;=\;10^{-\tau_{\text{dB}}/20}.$ For each TF bin $(f,u)$, the admissible magnitude change is bounded by $|\Delta_{f,u}|\;\leq\;\eta\,M_{f,u}$. We enforce these bounds via projection (clipping) after each optimizer step. Unlike loss-only quality terms (as in Li et al.), this explicit budget provides stronger and more reliable quality control and avoids re-embedding caused by overly aggressive updates. Bin-wise perceptual budgeting is common in classical, masking-inspired watermarking, but it has been largely de-emphasized in recent DL-based systems, where perceptual control is typically folded into a soft term in the overall loss that offers weaker guarantees than the hard constraints. In future work the tolerance $\tau_{\text{dB}}$ can be replaced by frequency-dependent thresholds or full psychoacoustic models (threshold-in-quiet and simultaneous masking), enabling budgets that also consider neighboring-frequency masking.
@@
Detector architecture. The detector architecture (illustrated in Figure  1) is designed so that activations across all layers remain stable under common audio edits. Robustness thus derives from the architecture itself rather than from optimization or distortion–simulation surrogates which tend to drive the system toward overfitting.
The first layer of the detector computes a Mel–spectrogram from the STFT magnitude. The Mel domain aggregates spectral energy into perceptually motivated bands, yielding a representation that is more robust than raw spectrograms to mild time–frequency distortions. Moreover, Mel features are standard in TTS/vocoder pipelines, increasing the chance that a watermark detectable in Mel-space survives voice cloning.
Following ablation findings in Li et al., we include a single pooling layer. This coarsens temporal resolution and improves robustness to local jitter and minor desynchronization.
The next stage contains three feature extraction blocks that intentionally avoid temporal mixing by treating Mel bands as channels and applying convolution along time dimension (kernel size =1=1, stride =1=1), followed by instance normalization and a Leaky ReLU activation. In this way, each time frame is processed independently, so the convolution kernels act as channel mixers across frequency at a fixed time, thereby avoiding sensitivity to cuts and re-ordering, that would otherwise induce activation-statistics drift.
Fully connected (FC) layers bind decisions to absolute positions and fixed input lengths. Deletions and splicing change the index–to–time mapping and lead to brittle, non-invariant activations. Moreover, FC layers require length standardization or padding that obscures real edit patterns. We therefore exclude FC layers entirely from the detector.
Instead, we introduce a Bitwise Readout Head (BRH) that reads out $N$ bits using paired convolutional filters. These filters aggregate evidence over time, and produce one position-agnostic score per bit. Concretely, the BRH applies two filter banks to the extracted features $Z\in\mathbb{R}^{C\times U^{\prime}}$:
@@
Global averaging in BRH removes the time axis before the final decision, ensuring that cropping, splicing, or frame deletions change the amount of evidence but not its required position. Because the feature extractor never relies on multi-frame receptive fields (kernel size for all convs is $1$), activation statistics remain stable under variable-length inputs and missing frames, thus enabling reliable fragment-level detection.
@@ 4 Experimental Setup
We compare against some of the strongest publicly available baselines (WavMark and AudioSeal) on VCTK [19] and LibriSpeech [20] datasets at $16$ kHz sampling rate. Perceptual quality is evaluated with PESQ [21] and STOI [22], while watermark robustness is measured by bit error rate (BER).
For a fair comparison, all methods are tested at a payload of $16$ bps, matching the training configuration of the comparative baselines, although AWARE exhibits similar performance at capacities exceeding $20$ bps. Adversarial embedding is optimized for $K=500$ iterations using the NAdam optimizer [23] (learning rate $0.1$), with a reduce-on-plateau scheduler (factor $0.9$). The push-loss margin weight is set to $\lambda=0.1$.
@@
[…]
5 Results and Analysis
Table 1 shows that AWARE achieves high audio quality and speech intelligibility, although slightly below AudioSeal, but above WavMark. Nevertheless, these scores indicate that the watermark remains imperceptible or near-imperceptible at all times. Given our goal of stronger cross-attack robustness, a modest quality compromise is expected.
@@ 5 Results and Analysis
WavMark $3.96\pm 0.43$ $0.96\pm 0.008$
AudioSeal $4.32\pm 0.29$ $0.99\pm 0.003$
AWARE $4.08\pm 0.37$ $0.97\pm 0.002$
Robustness results in Table 2 demonstrate that AWARE remains effective across diverse edits, with only small fluctuations in BER between conditions. In contrast, AudioSeal degrades sharply on spectral edits (e.g., LPF and BSF), while WavMark fails under coarse quantization and compression (PCM $8$-bit and MP3 $64$ kbps). Under pink-noise corruption, AWARE outperforms comparative methods, indicating strong resilience to background noise effects. Passage through a neural vocoder yields very low BER for AWARE, whereas baselines struggle. Temporal edits remain the hardest case. Under sample deletions and time-scale modifications AWARE incurs higher BERs, though still within a usable range.
Table 2: BER (%) under various edit/attack conditions.
Condition WavMark AudioSeal AWARE
Original 0.00 0.00 0.00
LPF 0.00 $14.58$ 0.00
HPF 0.00 $7.08$ 0.00
BSF 0.00 $33.81$ 0.95
PCM $24.46$ 1.47 1.43
MP3 $24.12$ 0.24 $0.71$
PN $28.59$ $10.89$ 1.61
RS 0.00 0.00 0.00
SD $1.43$ 0.70 $3.74$
TS $9.98$ $10.58$ 5.53
PS $50.00$ $2.55$ 0.92
NV $50.00$ $39.01$ 1.61

### v2

@@ Abstract
Prevailing practice in learning-based audio watermarking is to pursue robustness by expanding the set of simulated distortions during training. However, such surrogates are narrow and prone to overfitting. This paper presents AWARE (Audio Watermarking via Adversarial Resistance to Edits), an alternative approach that avoids reliance on attack-simulation stacks and handcrafted differentiable distortions. Embedding is obtained through adversarial optimization in the time–frequency domain under a level-proportional perceptual budget. Detection employs a time–order–agnostic detector with a Bitwise Readout Head (BRH) that aggregates temporal evidence into one score per watermark bit, enabling reliable watermark decoding even under desynchronization and temporal cuts. Empirically, AWARE attains high audio quality and speech intelligibility (PESQ/STOI) and consistently low BER across various audio edits, often surpassing representative state-of-the-art learning-based systems.
@@ 1 Introduction
Digital watermarking experienced its first major wave of research activity in the 1990s alongside the rapid proliferation of the Internet. Early systems were primarily designed for copyright protection and digital rights management (DRM), with the seminal work of Cox et al. introducing spread-spectrum principles to watermarking and setting the agenda for robustness-focused design Cox et al. (1997). Subsequent developments broadened the methodological toolbox with techniques such as quantization index modulation (QIM) Chen and Wornell (2001) and patchwork-style techniques Bender et al. (1996). While initial approaches were conceived as modality-agnostic and applicable across multimedia, the field soon bifurcated into image- and audio-specific lines of work Potdar et al. (2005); Hua et al. (2016), each exploiting modality characteristics to improve embedding efficiency and detection reliability. The dominant use case remained copyright protection, driven by the rise of large-scale online content distribution and associated piracy.
@@
Despite their relevance, traditional audio watermarking techniques still exhibit persistent limitations. Beyond insufficient robustness to classical signal processing operations, two challenges are particularly prominent: desynchronization and waveform cuts. Systems often include dedicated synchronization codes to address time-scale modifications, resampling drift, cropping, and jitter. However, reliably detecting synchronization markers is nearly as difficult as extracting the watermark itself, and thus inherits similar failure modes under distortion. Moreover, many legacy designs embed watermark bits within a single frame or a narrow group of frames, yielding limited temporal redundancy and weak fragment-level detectability. Missing or re-ordered frames, or partial content removal, can therefore break the decoding process and make the watermark hard or impossible to reconstruct.
@@
The advent of modern generative AI has precipitated a renaissance in digital audio watermarking. High-fidelity synthesis models such as: GANs Lee et al. (2022) and diffusion models Liu et al. (2023) enable convincing audio and audiovisual “deepfakes” at scale. The risks span reputational harm, fraud, misinformation, and weakened evidence reliability. In response, watermarking has re-emerged as a practical mechanism to label both synthetic and authentic content to support provenance, traceability, and downstream moderation. Accordingly, policy frameworks increasingly cite watermarking among key techniques for AI transparency and content provenance European Union (2024).
Concurrently, the community has begun to “fight fire with fire,” developing end-to-end deep learning (DL) audio watermarking systems. These methods have substantially improved robustness to many standard signal processing distortions and alleviated several long-standing limitations of classical watermarking. Leading this line of research, RobustDNN Pavlović et al. (2022) defined the basic blueprint, after which WavMark Chen et al. (2023) and AudioSeal Roman et al. (2024) introduced meaningful improvements. Nevertheless, contemporary benchmark studies still indicate unresolved limitations, with a clear room for progress Kovačević et al. (2025); Liu et al. (2024). Watermark decoding for audio under temporal cuts and splicing remains underexplored, aside from zero-bit approaches Pavlović et al. (2023) that have limited practical scope. The effects of transmission over the Web are also insufficiently studied, despite their practical importance, as such transmission typically introduces a compound set of distortions whose joint impact on watermark detection can be more severe than that of the individual degradations alone Pavlović et al. (2022).
Recent learning-based watermarking methods typically pursue robustness by training against a set of simulated distortions. While effective for the specific attacks seen during training, this strategy suffers from two fundamental limitations. First, the space of plausible audio edits is effectively unbounded, making exhaustive simulation infeasible. Second, models tend to overfit to the chosen distortion set and generalize poorly to unseen edits, especially when real-world degradations occur in combination and produce compound effects that are more harmful than those of the individual attacks considered separately.
In this work, we argue that robustness to such edits should primarily arise from architectural design choices rather than from increasingly complex attack-simulation pipelines. Specifically, audio watermark detection must aggregate weak, distributed evidence over time in a manner that is invariant to ordering, alignment, and signal length. Motivated by this perspective, we introduce AWARE (Audio Watermarking via Adversarial Resistance to Edits), a watermarking framework that combines adversarial embedding under explicit perceptual constraints with a detector architecture explicitly designed for robust audio watermark detection.
2 Threat model
We consider a practical audio watermarking threat model in which an adversary aims to disrupt watermark detection while keeping edits imperceptible or near-imperceptible to human listeners. Evaluations are restricted to black-box attacks, assuming no access to detector internals and allowing only signal-level manipulations. This setting reflects realistic deployments, where watermarking systems operate as closed modules and internal details are unavailable to attackers. Accordingly, we target common post-processing and distribution edits, including filtering, lossy compression, additive noise corruption, temporal cuts/cropping, desynchronization (e.g., resampling or time-scale modification), acoustic replay, networked transmission effects (e.g., codec distortion, delay, jitter, and packet loss). We also treat montage/splicing manipulations as in-scope, where content may be cut, removed, or re-ordered. Furthermore, we consider passage through speech-generation and transformation pipelines, including neural vocoder re-synthesis and neural audio compression.
Speech enhancement/separation Subakan et al. (2021) and voice cloning Wang et al. (2023) are also increasingly relevant. However, these pipelines introduce substantial signal and even semantic alterations. Robustness under such transformations remains an important open problem, as no established solution currently exists in the audio watermarking literature.
@@ 3 Background
In the standard adversarial setting, we start with a clean input $x$, label $y$, a model $f$, and a task-specific loss $\mathcal{L}$. The objective is to find a norm-bounded perturbation $\Delta$ that maximizes the loss: $\max_{\lVert\Delta\rVert_{p}\leq\epsilon}\ \mathcal{L}\big(f(x+\Delta),\,y\big)$. The choice of $p$ (e.g., $p=\infty$ or $2$) and $\varepsilon$ defines the perturbation “budget”.
Adversarial perturbations have proved effective across diverse tasks within the audio modality Yuan et al. (2018); Yu et al. (2021, 2023), predominantly as tools to break models rather than to enforce desired behaviors. Recent work has explored adversarial formulations of watermark embedding, where watermark is treated as a signal perturbation, optimized to steer the detector’s outputs. Li et al. Li et al. (2025) propose such watermarking framework for images. While the underlying idea of detector-driven adversarial embedding highlights useful principles for audio watermarking, it does not transfer directly to audio signals. In particular, audio watermark detection must contend with variable-length inputs, temporal re-ordering, and partial deletions, which fundamentally differ from the spatial cropping in images.
In our method, we adopt adversarial formulation for audio watermarking and avoid explicit simulation of attacks during optimization. Robustness is instead promoted through the design of a detector architecture tailored to the requirements of audio watermark detection, while watermark embedding is constrained by an explicit, fixed perceptual budget to control signal quality. These choices allow robustness to arise from architectural and optimization constraints rather than from exposure to specific simulated transformations.
@@ 4 AWARE: Method
Watermark Embedding. Embedding is carried out in the time-frequency (TF) domain, a standard setting for audio watermarking. We modify magnitudes under perceptual constraints and preserve the original phase for reconstruction (iSTFT). The complete procedure is outlined in Algorithm 1.
@@
Let $x\in\mathbb{R}^{T}$ be a waveform and let $\lvert\mathrm{STFT}(x)\rvert\in\mathbb{R}_{\geq 0}^{F\times U}$ denote its short-time Fourier transform (STFT) magnitude, with frequency bins $f\in[0,\dots,F-1]$ and time frames $u\in[0,\dots,U-1]$. We restrict perturbations to an audible midband $\mathcal{F}=\{f:f_{\ell}\leq f\leq f_{h}\}$ with $f_{\ell}=1000~\text{Hz}$ and $f_{h}=4000~\text{Hz}$ to avoid removal by low/high-pass filters.
Phase-domain watermark embedding is avoided due to human hearing being largely insensitive to changes in phase, which allows for low-audibility phase manipulations that effectively erase the mark.
We denote $w$ as a watermark of $N$ bits encoded antipodally as $\tilde{w}\in\{-1,+1\}^{N}$. This centers targets at zero, yields balanced gradients, and pairs naturally with margin-based objectives and sign decoding, unlike $\{0,1\}$ coding which pushes scores towards probability bounds and can hinder optimization. In our embedding procedure, the detector $D$ is randomly initialized and its weights are kept frozen throughout optimization. $D$ maps a TF magnitude representation of $x$ to $\left(-1,+1\right)^{N}$ and that provides gradients for the adversarial embedding process.
@@
Rather than imposing a single global norm budget on $\Delta$, we use a per-bin, level-proportional budget that allows larger changes where the signal is louder and smaller changes in quiet regions, consistent with basic psychoacoustics. Let $\tau_{\text{dB}}>0$ be a tolerance parameter (in dB). Its linear amplitude factor is $\eta\;=\;10^{-\tau_{\text{dB}}/20}.$ For each TF bin $(f,u)$, the admissible magnitude change is bounded by $|\Delta_{f,u}|\;\leq\;\eta\,M_{f,u}$. We enforce these bounds via projection (clipping) after each optimizer step. Unlike loss-only quality terms (as in Li et al. (2025)), this explicit budget provides stronger and more reliable quality control and avoids re-embedding caused by overly aggressive updates. Bin-wise perceptual budgeting is common in classical, masking-inspired watermarking, but it has been largely de-emphasized in recent learning-based systems, where perceptual control is typically folded into a soft term in the overall loss that offers weaker guarantees than the hard constraints.
In future work the tolerance $\tau_{\text{dB}}$ can be replaced by frequency-dependent thresholds or full psychoacoustic models (threshold-in-quiet and simultaneous masking), enabling budgets that also consider neighboring-frequency masking.
@@
Watermark detection. The detector (Fig. 1) is designed to keep activations across all layers stable under realistic audio edits. This leads to detection performance that is less dependent on the specific edit applied. Architectural depth and complexity are deliberately constrained. We argue that constraining architectural expressiveness improves robustness by discouraging reliance on fragile, edit-specific cues and by promoting aggregation of stable evidence. In particular, components that summarize information along the temporal axis reduce sensitivity to misalignment, partial content removal, and other realistic audio edits.
The first layer of the detector computes a Mel spectrogram from the STFT magnitude. The Mel domain aggregates spectral energy into perceptually motivated bands, yielding a representation that is more robust than raw spectrograms to mild time–frequency distortions. Moreover, Mel features are standard in TTS/vocoder pipelines, increasing the chance that a watermark detectable in Mel-space survives voice cloning.
Following ablation findings in Li et al. (2025), we include a single pooling layer. This coarsens temporal resolution and improves robustness to local jitter and minor desynchronization.
The next stage contains three feature extraction blocks that intentionally avoid temporal mixing by treating Mel bands as channels and applying 1D convolutions with kernel size $1$ along the temporal axis (stride $1$), followed by instance normalization and a Leaky ReLU. This processes each frame independently, so the filters act as channel (frequency) mixers at a fixed time and avoid temporal mixing.
Instead of using fully connected (FC) layers which bind decisions to absolute positions and fixed input lengths and become brittle under the effects that alter the index–time mapping, we introduce a Bitwise Readout Head (BRH) that reads out $N$ bits using paired filters. These filters aggregate evidence over time, and produce one position-agnostic score per bit. Concretely, the BRH applies two filter banks to the extracted features $Z\in\mathbb{R}^{C\times U^{\prime}}$:
@@
In addition to bit extraction, the BRH can be used to determine whether a given audio snippet is watermarked. Let $y=\mathrm{BRH}(x)$ denote the vector of per-bit scores produced by the detector. We define the confidence of the watermark decision as $\mathrm{conf}(x)\;=\;\frac{1}{N}\sum_{i=1}^{N}|y_{i}|$, i.e., the mean absolute activation of all BRH outputs. Intuitively, this score measures the overall strength of the aggregated watermark evidence. A detection threshold $\tau_{\text{conf}}$ is applied to determine watermark presence. If $\mathrm{conf}(x)>\tau_{\text{conf}}$, the signal is considered watermarked; otherwise, it is treated as non-watermarked. This is particularly useful under splicing and cutting, as BRH activations peak when the detector is aligned with segments that contain the watermark. Consequently, the confidence score can be evaluated over sliding or segmented windows and localize watermarked regions within a longer audio.
@@ 5 Experimental Setup
We compare against the strongest publicly available baselines (WavMark and AudioSeal) on $2000$ approximately $5$-second samples from the LibriSpeech Panayotov et al. (2015) dataset at a $16$ kHz sampling rate. Perceptual quality and speech intelligibility are evaluated using PESQ Rix et al. (2001) and STOI Taal et al. (2011), while watermark robustness is measured by bit error rate (BER). We evaluate message-level watermark detection using the false negative rate (FNR). The detection threshold $\tau_{\text{conf}}$ is set such that $\text{FPR}<0.01$ under no perturbations ($0.04$ in our setting).
For a fair comparison, all methods are tested with payloads of $16$ bps, matching the configuration of the comparative baselines. Detection thresholds are calibrated to achieve a comparable $\text{FPR}<0.01$ on clean, held-out audio. We note that the original AudioSeal evaluation tunes thresholds separately for each attack. We instead use a fixed threshold of $0.5$ across all conditions, as per-attack thresholding assumes prior knowledge of distortions and may not reflect realistic deployment settings.
In addition, we report results for AWARE at $20$ bps capacity by partitioning the signal into $1$ s segments and embedding $20$ bits per segment. Detection is performed by sliding-window with a fixed hop of $40$ ms and retaining windows whose BRH confidence exceeds a specified threshold $\tau_{\text{conf}}$, after which the final message is obtained by bitwise majority voting across detections. Segment length is adjusted to the smallest multiple of the hop size that is greater than or equal to the original length. This ensures full frame coverage in the STFT and prevents sample loss during the iSTFT reconstruction. For this higher-capacity setting, $\tau_{\text{conf}}$ is increased to $0.06$ to maintain a comparable false-positive rate.
Adversarial embedding is optimized for $K=500$ iterations using the NAdam optimizer Dozat (2016) (learning rate $0.1$), with a reduce-on-plateau scheduler (factor $0.9$). The push-loss margin weight is set to $\lambda=0.1$. STFTs are computed with a frame length of $1024$, hop length $256$, and a Hann window, with embedding restricted to [1000, 4000][1000,\,4000] Hz and detection using $128$ Mel bands. After optimization, the signal is reconstructed via iSTFT before detection.
Robustness is probed under: low/high-pass filtering (LPF at $4$ kHz, HPF at $1000$ Hz), PCM quantization to $8$ bits, MP3 compression at $64$ kbps, Gaussian and pink noise (GN; $20$ dB, PN; peak at $0.03$), resampling (RS) to $32$ kHz, and time–scale modification (TS; $\pm 35\%$). We evaluate resistance to band-stop (notch) filtering (BSF) with a $200$ Hz-wide notch. The notch center is swept across the 040004000 Hz range in $100$ Hz steps, and we report the worst-case error rate over all positions. Recent work reports vulnerabilities to such notches in models like AudioSeal Bas and Butora (2025). We test pitch shift (PS; $10$ cents), vocoder resynthesis (NV) using BigVGAN Lee et al. (2022), and neural audio compression (NAC) with EnCodec Défossez et al. (2022), for which benchmarks Kovačević et al. (2025); Liu et al. (2024) indicate limited robustness across many systems.
To explicitly evaluate montage and splicing scenarios, we consider two attacks. Segment removal (SR) removes a single contiguous segment corresponding to $50\%$ of the audio. Permutation attack (PA) simulates re-ordering by splitting the signal at a random point and swapping the segments.
@@
Figure 2: Quality–robustness trade-off controlled by the perceptual budget $\tau_{\text{dB}}$.
@@
WavMark $4.23\pm 0.18$ $0.99\pm 0.009$
AudioSeal $4.43\pm 0.08$ $0.99\pm 0.002$
AWARE $4.26\pm 0.09$ $0.99\pm 0.003$
AWARE (20bps) $4.24\pm 0.12$ $0.99\pm 0.004$
6 Results and Analysis
The quality–robustness trade-off (BER vs. PESQ) across distortions is shown in Fig. 2. The proposed model consistently operates in the upper-left region, corresponding to low BER and high perceptual quality. This indicates that the robustness comes with minimal perceptual cost and provides a practical region for parameter selection. Based on this trade-off, we select $\tau_{\mathrm{dB}}=6.2$ as the operating point for the remaining experiments reported in Tables 1 and 2. Table 1 shows that all methods achieve similarly high audio quality and speech intelligibility. Overall, the watermark remains imperceptible or near-imperceptible across systems.
Robustness results in Table 2 demonstrate that AWARE remains effective across diverse edits, with only small fluctuations in BER between conditions. AudioSeal degrades sharply on spectral edits (LPF, HPF and BSF), while WavMark fails under PCM and PS. Under noise corruption (GN and PN), AWARE outperforms comparative methods, indicating strong resilience to background noise and suggesting that the AWARE watermark is less noise-like and more structurally integrated into the signal. Passage through a NV yields very low error rate for AWARE, whereas baselines struggle. AudioSeal outperforms AWARE for NAC, likely due to its explicit exposure to such distortions during training. RS does not significantly impact any of the evaluated systems. Under TS, AWARE attains the lowest BER among all evaluated methods. SR proves to be a challenging case, as AWARE incurs higher BER, though still within a usable range and comparable to the strongest baseline. PA has little effect on AWARE, which confirms that time-order-agnostic aggregation mitigates sensitivity to reordering.
Table 2: BER with message-level FPR/FNR in parentheses.
Cond. WavMark AudioSeal AWARE AWARE (20 bps)
Original 0.000 (0.00/0.00) 0.000 (0.00/0.00) 0.000 (0.00/0.00) 0.000 (0.00/0.00)
LPF 0.000 (0.00/0.00) 0.146 (0.00/0.00) 0.000 (0.00/0.00) 0.000 (0.00/0.00)
HPF 0.000 (0.00/0.00) 0.409 (0.00/0.00) 0.000 (0.00/0.00) 0.000 (0.00/0.00)
BSF 0.000 (0.00/0.00) 0.338 (0.00/1.00) 0.010 (0.00/0.00) 0.000 (0.00/0.01)
PCM 0.100 (0.00/0.11) 0.005 (0.00/0.01) 0.000 (0.00/0.00) 0.000 (0.00/0.00)
MP3 0.023 (0.00/0.02) 0.000 (0.00/0.00) 0.001 (0.00/0.00) 0.003 (0.00/0.00)
PS 0.500 (0.00/0.75) 0.100 (0.00/0.28) 0.014 (0.00/0.00) 0.004 (0.00/0.00)
GN 0.460 (0.00/0.51) 0.033 (0.00/0.06) 0.012 (0.00/0.01) 0.000 (0.00/0.14)
PN 0.286 (0.00/0.75) 0.087 (0.00/0.68) 0.015 (0.00/0.02) 0.012 (0.00/0.09)
NV 0.500 (0.00/1.00) 0.390 (0.00/1.00) 0.016 (0.00/0.00) 0.009 (0.00/0.00)
NAC 0.500 (0.00/1.00) 0.091 (0.00/0.07) 0.179 (0.00/0.70) 0.260 (0.00/0.98)
RS 0.000 (0.00/0.00) 0.000 (0.00/0.00) 0.000 (0.00/0.00) 0.000 (0.00/0.00)
TS 0.257 (0.00/0.19) 0.068 (0.00/0.93) 0.016 (0.00/0.00) 0.027 (0.00/0.40)
SR 0.037 (0.00/0.07) 0.007 (0.00/0.21) 0.063 (0.00/0.10) 0.021 (0.00/0.04)
PA 0.000 (0.00/0.00) 0.042 (0.00/0.37) 0.000 (0.00/0.00) 0.000 (0.00/0.00)
RA 0.500 (0.00/1.00) 0.500 (0.00/1.00) 0.024 (0.00/0.11) 0.029 (0.00/0.58)
VT 0.239 (0.00/0.44) 0.480 (0.00/1.00) 0.102 (0.00/0.23) 0.104 (0.00/0.16)
The two compound, real-world attacks further highlight the benefits of the proposed design. Under RA, AWARE retains reliable detection, whereas both baselines fail. Similarly, under VT, AWARE significantly outperforms comparative methods. These results indicate that robustness achieved through architectural design generalizes beyond isolated signal-level edits to more realistic deployment scenarios.
AWARE ($20$ bps) follows similar trends and can outperform utterance-level embedding in several conditions (BSF, GN, PN, NV, and SR) due to sliding-window inference with majority voting, which introduces temporal redundancy and increase the chance of high-confidence detections. For scenarios with higher message-level FNR (NAC, TS, and RA), the number of available windows appears insufficient given the short LibriSpeech utterances ($\approx 5$ s). Longer signals should enable more reliable aggregation and improved detection. Overall, the results suggest that AWARE scales well with increased payload.